Set up an algebraic equation and solve the problem.
The ratio of male students to female students at a certain university is 3 to 19. If there is a total of 19,800 students, find the number of male students and the number of female students.
male students _______
female students _______

Answer by richwmiller(17219)

3/5=m/f,
3f/5=m
m+f=19800
3f/5+f=19800
8/5f=19800
2475*5=f
f = 12375, m = 7425